# Learn from Heterophily: Heterophilous Information-enhanced Graph Neural Network

### v1

@@ Abstract
Under circumstances of heterophily, where nodes with different labels tend to be connected based on semantic meanings, Graph Neural Networks (GNNs) often exhibit suboptimal performance. Current studies on graph heterophily mainly focus on aggregation calibration or neighbor extension and address the heterophily issue by utilizing node features or structural information to improve GNN representations. In this paper, we propose and demonstrate that the valuable semantic information inherent in heterophily can be utilized effectively in graph learning by investigating the distribution of neighbors for each individual node within the graph. The theoretical analysis is carried out to demonstrate the efficacy of the idea in enhancing graph learning. Based on this analysis, we propose HiGNN, an innovative approach that constructs an additional new graph structure, that integrates heterophilous information by leveraging node distribution to enhance connectivity between nodes that share similar semantic characteristics. We conduct empirical assessments on node classification tasks using both homophilous and heterophilous benchmark datasets and compare HiGNN to popular GNN baselines and SoTA methods, confirming the effectiveness in improving graph representations. In addition, by incorporating heterophilous information, we demonstrate a notable enhancement in existing GNN-based approaches, and the homophily degree across real-world datasets, thus affirming the efficacy of our approach.
Keywords:
@@
Graph Neural Networks Graph Homophily Heterophilous Information.
1 Introduction
@@
Graph-structured data widely exists in various applications including social networks[8] and recommendation systems[21] where entities are connected by their relations. To learn representations from these relational graphs, Graph Neural Networks(GNNs)[10, 17, 9] are proposed and have made substantial advancements. The effectiveness of GNNs on graph-structured data largely hinges on the homophily assumption, which posits that connected nodes tend to be similar[25]. Under this assumption, information is aggregated from neighboring nodes with similar features, facilitating effective information exchange within local structures. However, the homophily assumption does not always hold in many real-world scenarios where dissimilar nodes are connected[26], a situation referred to as heterophily or non-homophily. Heterophily poses a limitation to the performance of GNNs as it involves the aggregation of information from dissimilar nodes, thereby introducing additional noise during the aggregation process.
@@
To tackle the heterophily issue in graphs, several approaches have been proposed. Some methods calibrate the aggregation process by integrating high-order neighbors with different weights[1, 28, 4, 18], introducing adjacency matrices as a new feature of nodes[12], or propagating prior belief estimations[27]. While these aggregation calibration methods have demonstrated some improvements on heterophily graphs, they primarily focus on local neighbors and fail to aggregate information from global neighbors. Consequently, alternative methods are proposed to aggregate signals from global neighbors by integrating new filters[2, 13], or construct new graph structures[14, 23, 16, 11].
Despite the demonstrated effectiveness of these methods for heterophilous graphs, they overlook the information of heterophily during graph representation learning process, resulting in a suboptimal problem. Existing methods characterize the node information in graphs using features, structures, or aggregated neighbor features, yet none of them explain how nodes become homophilous or heterophilous.
@@
In this paper, we measure the heterophilous information as neighbor distribution, which describes the probability of a node’s neighbors belonging to specific classes. This information not only provides insight into the degree of homophily or heterophily but also enhances our understanding of the heterophilous contexts of nodes. We conduct both theoretical and empirical analyses to assess the impact of heterophilous information on graph learning. The results show that heterophilous information is effective in identifying semantically similar nodes. Building upon this analysis, we propose a novel method, HiGNN (Heterophilous Information-enhanced Graph Neural Network), designed to incorporate heterophilous information into GNNs. Leveraging this heterophilous information, the HiGNN demonstrates superior performance on both the homophilous and heterophilous datasets, significantly improving the homophily degree within newly constructed graph structures. In summary, our main contributions are summarized as follows:
@@
  • We first define the conception of heterophilous information in graphs and validate the effectiveness in improving homophily through both theoretical and empirical analysis.
  • We propose HiGNN, a novel method that leverages heterophilous information in GNNs. By constructing a new adjacency matrix based on the similarity of heterophilous information, HiGNN establishes new connections for nodes with similar heterophilous semantics, resulting in an improved homophily degree across different levels of homophily.
@@
  • Our experimental results demonstrate the superiority of our method on both the homophilous and heterophilous datasets. We also show improvements over general GNN baselines by incorporating heterophilous information.
2 Background
@@
We denote $\mathcal{G=(V,E)}$ as an undirected graph $\mathcal{G}$ with node set $\mathcal{V}$ and edge set $\mathcal{E}$. Let $A$ be an adjacency matrix where $A_{ij}=1$ or $A_{ij}=0$ represents the presence or absence of an edge $e_{ij}$ connecting node $v_{i}$ and $v_{j}$. For each node $v_{i}$, we use $\mathcal{N}_{i}=\{v_{j}|A_{ij}=1\}$ to denote its 1-hop neighbor set and $d_{i}$ to denote its node degree. We represent the node features as $X=\{x_{1},x_{2},...x_{n}\}\in\mathbb{R}^{N\times M}$ and the labels as $Y=\{y_{1},y_{2},...,y_{n}\}\in\mathbb{R}^{N\times 1}$.
@@
$h_{edge}=\frac{|\{{e_{ij}|e_{ij}\in\mathcal{E},y_{i}=y_{j}}\}|}{|\mathcal{E}|},h_{node}=\frac{1}{|\mathcal{V}|}\sum_{v_{i}\in\mathcal{V}}\frac{|\{{v_{j}|v_{j}\in\mathcal{N}_{i},y_{i}=y_{j}}\}}{d_{i}}$ (3)
@@
3 Methodology
In this section, we first introduce the definition of heterophilous information in Sect. 3.1. Next, we analyze how heterophilous information improves graph learning in Sect. 3.2. Finally, we introduce our proposed method HiGNN in Sect. 3.3.
3.1 Heterophilous Information
@@
3.2 Analysis of Heterophilous Information
After defining the new adjacency matrix $A^{\prime}$ with heterophilous information, we analyze how well this approach improves graph learning under general circumstances. Therefore, we assume each entry of $\mathcal{H}$ is corrupted by a Gaussian noise $\epsilon\sim N(0,\sigma^{2})$. Then we analyze the performance of graph learning with $A^{\prime}$ by theoretically examining the homophily degree $\hat{h}$ of $A^{\prime}$(proofs in Appendix A).
Theorem 3.1
The homophily degree $\hat{h}$ of $A^{\prime}$ is directly proportional to the threshold $\delta$, inversely proportional to the variance $\sigma^{2}$ of sampling noise, and directly proportional to the distance between the original homophily degree and reciprocal of c:
@@ 3.2 Analysis of Heterophilous Information
(a) Influence of $\sigma$ to $\hat{h}$
(b) Influence of $\delta$ to $\hat{h}$
(c) Influence of $c$ to $\hat{h}$
@@
To better illustrate the relationships, we plot the influences of these parameters. As depicted in Fig. 1a and Fig. 1b, $\hat{h}$ is directly proportional to $\delta$ and inversely proportional to $\sigma$ under different levels of $h$. These results align with our intuition: a larger noise during neighbor sampling makes it more challenging to capture their true neighbor distribution, thereby reducing $\hat{h}$. Moreover, a higher $\delta$ allows fewer edges with a higher similarity of neighbor distribution to connect, thereby increasing $\hat{h}$. Regarding the influence of $h$, as shown in Fig. 1c, an larger distance between $h$ and $\frac{1}{c}$ improves $\hat{h}$. When $h=\frac{1}{c}$, the probability of edges connecting to intra-class or inter-class nodes becomes identical i.e. $P(\boldsymbol{S^{+}}>\delta)=P(\boldsymbol{S^{-}}>\delta)$, which implies that there is no useful information from neighbors with respective to class. Conversely, if $h$ is very high or low, the neighbor distribution becomes distinguishable for each class, leading to an improved $\hat{h}$.
We then measure the influences of these parameters on real-world datasets. As depicted in Table 1, we measure the average standard deviation $\bar{\sigma}$, original homophily $h$, and the homophily improvement $\hat{h}-h$ under different threshold $\delta$ across 6 datasets. The results of $\hat{h}-h$ are shown in blue color if there is a positive improvement. The results align with our theoretical analysis: a larger $\delta$ improves $\hat{h}$ while a larger $\bar{\sigma}$ decreases $\hat{h}$. It is worth noting that the $\bar{\sigma}$ is typically small, resulting in a large improvement in $\hat{h}$ as shown in 1a. This shows the newly constructed $A^{\prime}$ effectively improves the graph learning. This is further evidenced in Table 1 where most datasets exhibit a significant increase of homophily with a high $\sigma$.
Table 1: Homophily Improvement on Heterophilous Datasets
@@ 3.3 HiGNN
Although this label estimation is not flawless and may result in a biased estimation of heterophilous information $\mathcal{H}$, the $\mathcal{H}$ defined in this paper is error-tolerant, accommodating differences between neighbor distributions in similar nodes. As illustrated in 1a, for $\delta$ that is close to $1$, the improved homophily $\hat{h}$ does not decrease significantly. Our experimental analysis of homophily improvement across 9 real-world datasets in Fig. 3 further substantiates this proposition.
@@ 3.4 Complexity Analysis
4 Experiments
4.1 Experimental Setup
Baselines. We benchmark HiGNN against a variety of methods, including non-graph methods, general GNNs, and heterophily-oriented GNNs. Specifically, we use MLP as the graph-agnostic method and GCN[10], GAT[17], SGC[19], Mixhop[1], and GraphSage[9] as the general GNNs baselines. For the heterophily-oriented GNNs, we use GCN2[3] H2GCN[28], LINKX[12], ACMGCN[13], and GPRGNN[4], GloGNN[11], WRGAT[16], and GGCN[23].
Datasets. We measure all the methods on 3 homophilous datasets [^1] https://github.com/kimiyoung/planetoid/tree/master/data and 6 heterophilous graph datasets [^2] https://github.com/CUAI/Non-Homophily-Large-Scale/tree/master/data. The statistics of all the datasets are presented in Table 2, where the homophilous graphs including Cora, CiteSeer, and PubMed, exhibit a significantly higher homophily compared to heterophilous graphs. We run 10 times for each dataset following public random splits with the ratio of 60%/20%/20% for train, validation, and test sets respectively [^3] https://github.com/bingzhewei/geom-gcn/tree/master/splits.
Table 2: Dataset Statistics
@@ 4.1 Experimental Setup
Implementations. To generate label predictions for neighbor distribution, we select ACM-GCN as an off-the-shelf model. Given the flexibility of our method, which can be integrated into any GNN method, we adopt GCN as the backbone model, enhanced with a high-pass filter[13]. Then we run HiGNN and all the baselines on 9 datasets, implemented using Pytorch on an A5000 GPU. The models are trained until convergence is achieved, with an early stopping criterion of 40 epochs or a maximum of 2000 epochs. For all the baselines, we search for the optimal hyper-parameters, including hidden embedding size, learning rate, weight decay, and dropout rate on the validation sets. Specifically for HiGNN, we search for the threshold $\delta$ from $[0.5,0.8,0.9,1.0]$ and lambda $\lambda$ from $[1,0.5,0.1,0.05,0.01]$ for each dataset.
4.2 Performance

### v2

@@ Abstract
Under circumstances of heterophily, where nodes with different labels tend to be connected based on semantic meanings, Graph Neural Networks (GNNs) often exhibit suboptimal performance. Current studies on graph heterophily mainly focus on aggregation calibration or neighbor extension and address the heterophily issue by utilizing node features or structural information to improve GNN representations. In this paper, we propose and demonstrate that the valuable semantic information inherent in heterophily can be utilized effectively in graph learning by investigating the distribution of neighbors for each individual node within the graph. The theoretical analysis is carried out to demonstrate the efficacy of the idea in enhancing graph learning. Based on this analysis, we propose HiGNN, an innovative approach that constructs an additional new graph structure, that integrates heterophilous information by leveraging neighbor distribution to enhance connectivity between nodes that share similar semantic characteristics. We conduct empirical assessments on node classification tasks using both homophilous and heterophilous benchmark datasets and compare HiGNN to popular GNN baselines and SoTA methods, confirming the effectiveness in improving graph representations. In addition, by incorporating heterophilous information, we demonstrate a notable enhancement in classicial GNN baselines across real-world datasets, thus affirming the efficacy of our approach.
Index Terms:
@@
Graph Neural Networks, Graph Homophily, Heterophilous Information, Graph Representation Learning.
I Introduction
@@
Graph-structured data widely exists in various applications including social networks[1] and recommendation systems[2] where entities are connected by their relations. To learn representations from these relational graphs, Graph Neural Networks(GNNs)[3, 4, 5, 6] are proposed and have made substantial advancements. The effectiveness of GNNs on graph-structured data largely hinges on the homophily assumption, which posits that connected nodes tend to be similar[7, 8, 9, 10, 11]. Under this assumption, information is aggregated from neighboring nodes with similar features, facilitating effective information exchange within local structures. However, the homophily assumption does not always hold in many real-world scenarios where dissimilar nodes are connected[12], a situation referred to as heterophily or non-homophily. Heterophily poses a limitation to the performance of GNNs as it involves the aggregation of information from dissimilar nodes, thereby introducing additional noise during the aggregation process.
@@
To handle the heterophily in graphs, several approaches have been proposed. Some methods calibrate the aggregation process by integrating high-order neighbors with different weights[13, 14, 15, 16], introducing adjacency matrices as a new feature of nodes[17], or propagating prior belief estimations[18]. While these aggregation calibration methods have demonstrated some improvements on heterophily graphs, they primarily focus on local neighbors and fail to aggregate information from global neighbors. Consequently, alternative methods are proposed to aggregate signals from global neighbors by integrating new filters[19, 20], or construct new graph structures[21, 22, 23, 24].
Figure 1: An academic collaboration graph with many interdisciplinary collaborations, such as scholars from Bioinformatics(BioInfo) and Financial Technology(FinTech), resulting in high heterophily with rich semantic information.
Although the heterophily is conventionally considered as noises in many graph-based machine learning tasks during message aggregation[20, 25, 26], there are scenarios where a heterophilous graph structure could carry meaningful semantic information. For example, as an academic collaboration network shown in Figure 1, scholars from different disciplines are connected by academic collaborations. The heterophily in this network reflects meaningful interdisciplinary collaborations, such as Bioinformatics (collaborations between Biology and Data Science) and Financial Technology (collaborations between Business and Data Science). In this case, a high degree of heterophily indicates a rich exchange of ideas and methodologies between different fields, fostering innovation and broadening research perspectives. While these heterophilous graphs contain rich semantic information, the performance of GNNs is limited by the heterophily. This leads to a crucial question: How can we utilize useful semantic information from the heterophily?
@@
To extract the semantic information from the heterophily, we first define the conception of heterophilous information, which describes the probability of a node’s neighbors belonging to specific classes. This information not only provides insight into the degree of homophily or heterophily but also enhances our understanding of the heterophilous contexts of nodes. As the previous example shown in Figure 1, scholars in Bioinformatics or Financial Technology share a high degree of heterophily with their neighbors. Although these heterophilous contexts are conventionally regarded as noise during message aggregation in GNNs, their heterophilous information can identify them as interdisciplinary collaborators in specific research areas. To better understand how effective the heterophilous information is in improving GNN performance, we conduct theoretical and empirical analyses to assess its impact. The results show that heterophilous information is effective in identifying semantically similar nodes. Building upon this analysis, we propose a novel method, HiGNN (Heterophilous Information-enhanced Graph Neural Network), designed to incorporate heterophilous information into GNNs. Leveraging this heterophilous information, the HiGNN demonstrates superior performance on both homophilous and heterophilous datasets, significantly improving the homophily degree within newly constructed graph structures. In summary, our main contributions are summarized as follows:
@@
  • The concept of heterophilous information in graphs is defined for the first time, and its effectiveness in improving homophily is verified through theoretical and empirical analysis.
  • The HiGNN method is proposed, which constructs a new adjacency matrix and establishes new connections between nodes based on the similarity of heterophilous information, thus distinguishing semantic information from noises and improving the performance of graph representation at different levels of heterophily.
@@
  • The experimental results demonstrate the superiority of our method on both the homophilous and heterophilous datasets. We also show improvements over existing GNN baselines by introducing heterophilous information.
II Background
@@
We denote $\mathcal{G=(V,E)}$ as an undirected graph $\mathcal{G}$ with node set $\mathcal{V}$ and edge set $\mathcal{E}$. Let $A$ be an adjacency matrix where $A_{ij}=1$ or $A_{ij}=0$ represents the presence or absence of an edge $e_{ij}$ connecting node $v_{i}$ and $v_{j}$. For each node $v_{i}$, we use $\mathcal{N}_{i}=\{v_{j}|A_{ij}=1\}$ to denote its 1-hop neighbor set and $d_{i}$ to denote its node degree. We represent the node features as $X=\{x_{1},x_{2},...x_{n}\}\in\mathbb{R}^{N\times M}$ and the labels as $Y=\{y_{1},y_{2},...,y_{n}\}\in\mathbb{R}^{N\times 1}$, where $N$ and $M$ refer to the number of nodes and features in $\mathcal{G}$ respectively.
@@
$\begin{split}h_{edge}&=\frac{|\{{e_{ij}|e_{ij}\in\mathcal{E},y_{i}=y_{j}}\}|}{|\mathcal{E}|}\\ h_{node}&=\frac{1}{|\mathcal{V}|}\sum_{v_{i}\in\mathcal{V}}\frac{|\{{v_{j}|v_{j}\in\mathcal{N}_{i},y_{i}=y_{j}}\}}{d_{i}}\end{split}$ (3)
@@
III Methodology
In this section, we first introduce the definition of heterophilous information in Sect. III-A. Next, we analyze how heterophilous information improves graph learning in Sect. III-B. Finally, we introduce our proposed method HiGNN in Sect. III-C.
III-A Heterophilous Information
@@
III-B Analysis of Heterophilous Information
After defining the new adjacency matrix $A^{\prime}$ with heterophilous information, we analyze how well this approach improves graph learning by theoretically examining the homophily degree $\hat{h}$ of $A^{\prime}$. Therefore, we first make the following assumptions:
  • The number of nodes is balanced across each class.
Then we have the following theorem under the assumptions.
Theorem III.1
The homophily degree $\hat{h}$ of $A^{\prime}$ is directly proportional to the threshold $\delta$, inversely proportional to the variance $\sigma^{2}$ of sampling noise, and directly proportional to the distance between the original homophily degree and reciprocal of the class number $c$:
@@ III-B Analysis of Heterophilous Information
$Proof.$ Following the aforementioned assumptions, we ensure the original homophily degree is $h$. Then we can get the heterophilous information of node $u$ as a vector:
$\mathcal{H}_{u}=[p_{1},p_{2},\cdots,p_{c}],\text{where}\ p_{i}=\begin{cases}h+\epsilon_{u,i}&\text{if }i=y_{u}\\ \frac{1-h}{c-1}+\epsilon_{u,i}&\text{if }i\neq y_{u}\end{cases}$ (7)
$\begin{split}\boldsymbol{S^{+}}&=\cos(\mathcal{H}_{u},\mathcal{H}_{v})\\ &=\frac{1}{\mathcal{H}_{norm}}\biggl{(}(h+\epsilon_{v,y_{v}})(h+\epsilon_{u,y_{u}})\\ &+\sum_{\begin{subarray}{c}i=1,\\ i\neq y_{u}\end{subarray}}^{c}(\frac{1-h}{c-1}+\epsilon_{u,i})(\frac{1-h}{c-1}+\epsilon_{v,i})\biggl{)}\\ \boldsymbol{S^{-}}&=\cos(\mathcal{H}_{u},\mathcal{H}_{v})\\ &=\frac{1}{\mathcal{H}_{norm}}\biggl{(}(h+\epsilon_{u,y_{u}})(\frac{1-h}{c-1}+\epsilon_{v,y_{v}})\\ &+(h+\epsilon_{v,y_{v}})(\frac{1-h}{c-1}+\epsilon_{u,y_{u}})\biggl{.}\\ \biggl{.}&+\sum_{\begin{subarray}{c}i=1,\\ i\neq y_{u},i\neq y_{v}\end{subarray}}^{c}(\frac{1-h}{c-1}+\epsilon_{u,i})(\frac{1-h}{c-1}+\epsilon_{v,i})\biggl{)}\end{split}$ (8)
[…]
where
$\begin{split}\epsilon^{+}&=h(\epsilon_{u,y_{u}}+\epsilon_{v,y_{v}})+\sum_{\begin{subarray}{c}i=1,i\neq y_{u}\end{subarray}}^{c}\frac{1-h}{c-1}(\epsilon_{u,i}+\epsilon_{v,i})\\ &\sim N(0,(h^{2}+\frac{(1-h)^{2}}{c-1})2\sigma^{2})\end{split}$ (11)
Then we can get the probability of homophilous edges as
Let
$t_{+}=\frac{h^{2}+\frac{(1-h)^{2}}{c-1}-\delta\mathcal{H}_{norm}}{(h^{2}+\frac{(1-h)^{2}}{c-1})^{\frac{1}{2}}\sqrt{2}\sigma}$ (13)
we have $P(\boldsymbol{S^{+}}>\delta)=\Phi(t_{+})$.
Similarly, for $\boldsymbol{S^{-}}$, we follow the same process as $\boldsymbol{S^{+}}$ and we can get $P(\boldsymbol{S^{-}}>\delta)=\Phi(t_{-})$, where
$t_{-}=\frac{2h\frac{1-h}{c-1}+(C-2)(\frac{1-h}{c-1})^{2}-\delta\mathcal{H}_{norm}}{(h^{2}+\frac{(1-h)^{2}}{c-1})^{\frac{1}{2}}\sqrt{2}\sigma}$ (14)
From the partial derivative results, we can get Theorem III.1 proved.
@@
To better illustrate how the improved homophily $\hat{h}$ is affected, we plot the influences of these parameters according to the theoretical results in Eq. (15). As depicted in Fig. 2a and Fig. 2b, $\hat{h}$ is directly proportional to $\delta$ and inversely proportional to $\sigma$ under different levels of $h$. These results align with our intuition: a larger noise during neighbor sampling makes it more challenging to capture their true neighbor distribution, thereby reducing $\hat{h}$. Moreover, a higher $\delta$ allows fewer edges with a higher similarity of neighbor distribution to connect, thereby increasing $\hat{h}$. For the influence of $h$, as shown in Fig. 2c, we can see that a larger distance between $h$ and $\frac{1}{c}$ improves $\hat{h}$. When $h=\frac{1}{c}$, the probability of edges connecting to intra-class or inter-class nodes becomes identical i.e. $P(\boldsymbol{S^{+}}>\delta)=P(\boldsymbol{S^{-}}>\delta)$, which implies that there is no useful information from neighbors because the aggregated messages from neighbors become indistinguishable for each class. Conversely, if $h$ is very high or low, the neighbor distribution becomes distinguishable for each class, leading to an improved $\hat{h}$. This phenomenon has also been observed in the theoretical analysis by [27], which is known as the mid-homophily pitfall.
We then measure the influences of these parameters on real-world datasets. As depicted in Table I, we measure the average standard deviation $\bar{\sigma}$, original homophily $h$, and the homophily improvement $\hat{h}-h$ under different threshold $\delta$ across 6 datasets. The results of $\hat{h}-h$ are highlighted in bold if there is a positive improvement. The results align with our theoretical analysis: a larger $\delta$ improves $\hat{h}$ while a larger $\bar{\sigma}$ decreases $\hat{h}$. It is worth noting that the $\bar{\sigma}$ is typically small, resulting in a large improvement in $\hat{h}$ as shown in 2a. This shows the newly constructed $A^{\prime}$ effectively improves the graph learning. This is further evidenced in Table I where most datasets exhibit a significant increase of homophily with a high $\sigma$. This observation validates the efficacy of leveraging heterophilous information to improve the performance of GNNs.
TABLE I: Homophily Improvement on Heterophilous Datasets
@@ III-C HiGNN
Although this label estimation is not flawless and may result in a biased estimation of heterophilous information $\mathcal{H}$, the $\mathcal{H}$ defined in this paper is error-tolerant, accommodating differences between neighbor distributions in similar nodes. As illustrated in 2a, for $\delta$ that is close to $1$, the improved homophily $\hat{h}$ does not decrease significantly. Our experimental analysis of homophily improvement across 9 real-world datasets in Fig. 4 further substantiates this proposition.
@@ III-D Complexity Analysis
IV Experiments
IV-A Experimental Setup
Baselines. We benchmark HiGNN against a variety of methods, including non-graph methods, general GNNs, and heterophily-oriented GNNs. Specifically, we use MLP as the graph-agnostic method and GCN[3], GAT[4], SGC[30], Mixhop[13], and GraphSage[5] as the general GNNs baselines. For the heterophily-oriented GNNs, we use GCN2[31], H2GCN[14], LINKX[17], ACMGCN[20], and GPRGNN[15], GloGNN[24], WRGAT[23], and GGCN[22].
Datasets. We measure all the methods on 3 homophilous datasets [^1] https://github.com/kimiyoung/planetoid/tree/master/data and 6 heterophilous graph datasets [^2] https://github.com/CUAI/Non-Homophily-Large-Scale/tree/master/data. The statistics of all the datasets are presented in Table II, where the homophilous graphs including Cora, CiteSeer, and PubMed, exhibit a significantly higher homophily compared to heterophilous graphs. We run 10 times for each dataset following public random splits with the ratio of 60%/20%/20% for train, validation, and test sets respectively [^3] https://github.com/bingzhewei/geom-gcn/tree/master/splits.
TABLE II: Dataset Statistics
@@ IV-A Experimental Setup
Implementations. To generate label predictions for neighbor distribution, we select ACM-GCN as an off-the-shelf model. Given the flexibility of our method, which can be integrated into any GNN method, we adopt GCN as the backbone model, enhanced with a high-pass filter[20]. Then we run HiGNN and all the baselines on 9 datasets, implemented using Pytorch on an A5000 GPU. The models are trained until convergence is achieved, with an early stopping criterion of 40 epochs or a maximum of 2000 epochs. For all the methods, we search for the optimal hyper-parameters, including hidden embedding size from $\{64,128,256\}$, learning rate from $\{10^{-2},10^{-3}\}$, weight decay from $\{5\times 10^{-4},5\times 10^{-5}\}$, number of layers from $\{1,2\}$ and dropout rate from $\{0.3,0.4,0.5,0.6,0.7\}$ on the validation sets for each dataset. Specifically for HiGNN, we search for the threshold $\delta$ from $\{0.5,0.8,0.9,1.0\}$ and lambda $\lambda$ from $\{1,0.5,10^{-1},5\times 10^{-2},1\times 10^{-2},1\times 10^{-3},1\times 10^{-4}\}$ to control the amount of heterophilous information involved. The final hyper-parameters for HiGNN are provided in Table V.
IV-B Performance